A personality test has a subsection designed to assess the honesty of the test-taker. Suppose that youre interested in the
Answer #1

Given ME = 3, \sigma = 22, \alpha = 0.05

The Zcritical at \alpha = 0.05 is 1.96

The ME is given by :

ME = Zcritical *

Squaring both sides we get: (ME)2 = (Z critical)2 * \sigma 2/n

Therefore n = (Zcritical * \sigma /ME)2 = (1.96 * 22/ 3)2 = 206.59

Therefore n = 207 (Taking it to the next whole number)

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

ME = Zcritical *

We were unable to transcribe this image

We were unable to transcribe this image